# Do LLMs Adhere to Label Definitions? Examining Their Receptivity to External Label Definitions

### v1

@@ 1 Introduction
This example reveals a fundamental challenge in how LLMs process conflicting information sources. When external definitions clash with a model’s training data, we see unpredictable behavior—sometimes the model follows the external guidance, other times it relies on internal knowledge (Burns et al., 2022; Duan et al., 2024). To understand this systematically, we need to examine two key aspects: how models handle Knowledge Conflict (when external definitions contradict training) and Definition Integration (how definitions are presented in prompts) (Azaria and Mitchell, 2023; Gaur et al., 2022). The LLaMA-3 case demonstrates both issues—the model’s internal understanding of “neutral” conflicts with the manipulated definition, yet it still attempts to integrate the external guidance into its reasoning. This suggests that explanation-based classification may be more vulnerable to definitional inconsistencies than previously assumed.
To better understand the role of definitions in explanation-based classification tasks using LLMs, we investigate the following questions: Q1: Do LLMs rely more on their internalized, potentially opaque representations, raising questions about the interplay between external guidance and inherent knowledge? Q2: Do they consistently adhere to external definitions? In this work, we restrict our study to classification tasks, as they provide clear ground truth labels that make it easier to assess whether models follow external definitions or revert to internal knowledge.
To answer these questions systematically, we conduct experiments across four diverse datasets: e-snli for general natural language inference (Camburu et al., 2018), wellxplain for mental health (Garg, 2024), hateexplain for hate speech detection (Mathew et al., 2021), and wice for fact-checking (Kamoi et al., 2023). This multi-domain approach allows us to test whether our findings generalize beyond any single task or domain. Our main contributions are as follows: 1) We develop evaluation strategies to assess the performance and explainability of LLMs when provided with label definitions, which serve as a lightweight and interpretable form of external knowledge injection. 2) We design two probing strategies, definition permutation and definitional perturbation, to assess the sensitivity of LLMs to definition quality and alignment. 3) We conduct extensive experiments across four diverse benchmarks, e-snli, wellxplain, hatexplain, and wice, to assess model receptivity to label definitions.
@@ 3 Setup and Methodology
To investigate definition receptivity across diverse model architectures, we selected four state-of-the-art LLMs with deliberately varied characteristics: GPT-4, LLaMA-3, Phi-3, and Mistral.
This selection spans from compact 3.8B parameter models to trillion-parameter architectures, enabling us to determine whether investments in larger models yield proportional improvements in definition adherence—a key consideration for AI deployment in high-stakes domains where consistent interpretation of instructions is essential. Since definitions are concise, their inclusion adds minimal overhead in terms of token length. Nonetheless, this setup is primarily designed for short to medium-length inputs. For tasks requiring longer contexts (e.g., document-level inputs), the approach may not be suitable for smaller models like Phi-3-mini-4k, which has a 4k token limit, thereby restricting the ability to integrate external definitions without truncation.
Each model represents distinct design philosophies with practical implications for deploying LLMs in definition-sensitive contexts: GPT-4 serves as our high-performance benchmark, representing proprietary models that organizations might access through APIs (Achiam et al., 2023); LLaMA-3 exemplifies open-source alternatives increasingly adopted in industry settings requiring customization and privacy (Meta, 2024); Phi-3-mini addresses the growing need for efficient, deployable models in resource-constrained environments (Bilenko, 2024); and Mistral-7B represents models optimized for the performance-efficiency trade-off crucial for commercial applications (Jiang et al., 2023). This diversity allows us to provide actionable insights on which architectural approaches best ensure definitional consistency across different deployment scenarios—whether in cloud infrastructure, edge devices, or hybrid settings. Detailed specifications for each model are provided in Table 5 (appendix˜A), with a representative prompt template shown in Table 6 (appendix˜A). For GPT-4, we balanced representation and cost by sampling 500 instances per label across datasets [^3] This approach ensures balanced class representation while managing computational resources—a practical constraint in production settings as well.. For all other models, we used complete test sets.
@@ 4 Experiments, Results, and Analysis
Model Permutation e-snli wellxplain hatexplain wice GPT-4 $\text{MCC}_{a}$ 0.83 0.35 0.42 0.32 $\overline{\text{MCC}}_{m}$ 0.74 0.17 0.27 0.16 LLaMA-3 $\text{MCC}_{a}$ 0.42 0.20 0.21 0.13 $\overline{\text{MCC}}_{m}$ 0.34 0.15 0.17 0.09 Phi-3 $\text{MCC}_{a}$ 0.24 0.32 0.20 0.05 $\overline{\text{MCC}}_{m}$ 0.23 0.11 0.06 0.04 Mistral $\text{MCC}_{a}$ 0.51 0.10 0.05 0.05 $\overline{\text{MCC}}_{m}$ 0.48 0.07 0.05 0.03
@@
e-snli wellxplain hatexplain wice Model $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ GPT-4 Meta-Response 0.83 0.83 0.31 0.31 0.38 0.31 0.36 0.44 0.13 0.31 0.32 LLaMA-3 0.08 0.32 0.36 0.16 0.20 0.21 0.10 0.13 0.15 0.09 0.13 0.13 Phi-3 0.22 0.21 0.23 0.29 0.35 0.38 0.11 0.08 0.14 0.04 0.01 0.05 Mistral 0.20 0.48 0.40 0.18 0.20 0.20 0.05 0.07 0.09 0.03 0.03 0.05
Table 2: Impact of Definition Quality: Incorrect, Slightly Incorrect, and Correct Definitions across the general domain dataset (e-snli) and domain-specific datasets (wellxplain and hatexplain). Results highlight distinct patterns in model performance, with LLaMA-3 showing steady improvement from Incorrect to Correct, while Mistral exhibits strong sensitivity to definition accuracy, particularly in domain-specific settings. Subscripts Inc., SInc., and Cor. denote Incorrect, Slightly Incorrect, and Correct, respectively. † This response reflects the model declining to answer due to conflicting definitions.
@@ 4.1 LLM Receptivity: Results and Analysis
In the definition permutation scenario, Table 1 shows clear performance differences between aligned and misaligned label-definition mappings. Across all datasets, models achieve significantly higher MCC scores with correctly permuted definitions (aligned), with Mistral on hatexplain being the striking exception where alignment produced minimal improvement—a surprising deviation from the otherwise consistent pattern. These findings align with the distributional hypothesis of language learning (Firth, 1957; Saxena et al., 2024), whereby LLMs learn word meanings through their contextual associations. When faced with conflicting cues—explicit definitions versus internalized patterns—models generally prioritize the explicit definitions provided in the prompt context, suggesting that instruction-following capabilities can override pre-trained associations.
The label definition accuracy scenario, summarized in Table 2, reveals varying sensitivities to definition quality across models and tasks. For e-snli, LLaMA-3 demonstrates remarkably strong responsiveness, with MCC values increasing more than threefold when moving from incorrect to slightly incorrect definitions, and nearly fourfold when comparing incorrect to correct definitions. Mistral shows substantial but less dramatic improvements (140% and 100% respectively), while Phi-3 exhibits a surprisingly minimal sensitivity with near-negligible changes (-4.55% and +4.55%)—an unexpected resilience to definitional variations. Most unexpectedly, GPT-4 surprised us entirely by detecting definitional inconsistencies and declining to provide predictions—a sophisticated meta-cognitive response none of the other models exhibited. This pattern reveals a key distinction between linguistic competence (internalized knowledge of language structures) and performance (how that knowledge is applied in specific contexts), with larger models demonstrating stronger metacognitive capabilities to detect contradictions between provided definitions and their pre-training.
@@
For the wellxplain mental health categorization task, all models show moderate improvements with better definitions, though the sensitivity varies. GPT-4 and LLaMA-3 show similar improvement patterns (22.58% and 25% respectively) when moving from incorrect to more accurate definitions. Phi-3 demonstrates greater definitional sensitivity, with performance improving by approximately one-fifth when using slightly incorrect definitions and by nearly one-third with correct definitions. Contrary to expectations, Mistral shows the least variation (11.11%), suggesting it relies more heavily on internal knowledge for this domain—a surprising finding given its stronger responsiveness in the e-SNLI task. This aligns with theories of knowledge acquisition, suggesting that when LLMs lack extensive pre-training exposure to specialized domains, they become more dependent on explicit contextual definitions, similar to how humans learn novel concepts by relying on explicit instruction when they lack prior experience.
In the hatexplain content moderation task, definition quality impacts performance substantially across all models. GPT-4’s accuracy improves steadily as definitions become more precise (16.13% with slightly incorrect and 41.94% with correct definitions). LLaMA-3 follows a similar pattern with more pronounced gains (30% and 50%). Interestingly, both Phi-3 and Mistral show unexpectedly large improvements when given correct definitions (75% and 80% respectively)—a stark contrast to their behavior on other tasks. This suggests a “knowledge integration threshold” where models shift from relying on internal representations to trusting external definitions when the latter are sufficiently precise and consistent with their training.
For the wice fact-checking task, model performance under varying definition accuracies reveals minimal responsiveness across all models. Unlike other datasets, the MCC scores remain nearly flat, with negligible gains when moving from incorrect to correct definitions. LLaMA-3 and Mistral show no improvement at all (0.09 → 0.13 and 0.03 → 0.05, respectively), while Phi-3 even declines slightly from slightly incorrect to correct. These results suggest that wice, as an evidence-retrieval task, is less sensitive to semantic label definitions and more reliant on factual pattern matching—highlighting a limitation in current LLMs’ ability to incorporate abstract definitional guidance into decision-making for fact-based inference tasks.
@@
e-snli wellxplain hatexplain wice Model Scenario MCCMCC RougeRouge MCCMCC RougeRouge MCCMCC RougeRouge MCCMCC $F1$ Vanilla 0.86 0.21 0.41 0.06 0.33 0.05 0.28 0.71 GPT-4 Fixed Definition 0.80 0.20 0.46 0.05 0.36 0.04 0.32 0.73 Adjusted Definition 0.83 0.20 0.55 0.07 0.33 0.06 0.35 0.69 Fixed Definition+Few-shot 0.70 0.26 0.46 0.10 0.30 0.31 0.32 0.73 Vanilla 0.47 0.18 0.29 0.04 0.14 0.31 0.14 0.69 LLaMA-3 Fixed Definition 0.35 0.15 0.16 0.04 0.23 0.18 0.13 0.69 Adjusted Definition 0.36 0.14 0.44 0.40 0.14 0.10 0.14 0.69 Fixed Definition+Few-shot 0.36 0.20 0.38 0.38 0.15 0.04 0.11 0.74 Vanilla 0.48 0.19 0.33 0.06 0.13 0.01 0.03 0.68 Phi-3 Fixed Definition 0.38 0.12 0.38 0.06 0.19 0.02 0.05 0.68 Adjusted Definition 0.46 0.15 0.49 0.14 0.17 0.01 0.12 0.67 Fixed Definition+Few-shot 0.31 0.22 0.62 0.13 0.16 0.10 0.02 0.69 Vanilla 0.57 0.17 0.20 0.07 0.01 0.47 0.03 0.70 Mistral Fixed Definition 0.32 0.20 0.20 0.07 0.12 0.35 0.05 0.69 Adjusted Definition 0.34 0.19 0.42 0.39 0.14 0.01 0.04 0.69 Fixed Definition+Few-shot 0.49 0.24 0.39 0.36 0.14 0.05 0.01 0.70
Table 3: Impact of Definitions Integration on Classification Performance and Explanation Alignment: (i) vanilla, (ii) fixed definition, (iii) adjusted definition, and (iv) definition + few-shot. All models achieve high performance in the vanilla scenario but exhibit superior explanation quality in the fixed definition+few-shot scenario for the e-snli task.
@@ 4.2 Definitions Integration: Results and Analysis
For e-snli, models perform best with the definition-free vanilla setting—a counterintuitive finding that challenges conventional wisdom about explicit guidance. While GPT-4 experiences modest degradation with definitions (6.98% and 3.49% decreases), LLaMA-3 and Mistral show more substantial performance drops (23.40% to 43.86%). This pattern aligns with the bayesian inference framework of in-context learning proposed by Xie et al. (2021), where models already possess robust internal representations of common linguistic tasks through pretraining and may experience interference when explicit definitions contradict these learned representations (Min et al., 2022). Most surprisingly, adding few-shot examples alongside definitions further reduces performance across all models, with GPT-4 dropping by 18.60%—suggesting that exemplars can actually create confusion when combined with abstract definitions for general language tasks.

### v2

@@ 1 Introduction
This example reveals a fundamental challenge in how LLMs process conflicting information sources. When external definitions clash with a model’s training data, we see unpredictable behavior; sometimes the model follows the external guidance, other times it relies on internal knowledge (Burns et al., 2022; Duan et al., 2024). To understand this systematically, we need to examine two key aspects: how models handle Knowledge Conflict (when external definitions contradict training) and Definition Integration (how definitions are presented in prompts) (Azaria and Mitchell, 2023; Gaur et al., 2022). The LLaMA-3 case demonstrates both issues—the model’s internal understanding of “neutral” conflicts with the manipulated definition, yet it still attempts to integrate the external guidance into its reasoning. This suggests that explanation-based classification may be more vulnerable to definitional inconsistencies than previously assumed.
To better understand the role of definitions in explanation-based classification tasks using LLMs, we investigate the following questions: Q1: Do LLMs rely more on their internalized, potentially opaque representations, raising questions about the interplay between external guidance and inherent knowledge? Q2: Do they consistently adhere to external definitions? In this work, we restrict our study to classification tasks, as they provide clear ground truth labels that make it easier to assess whether models follow external definitions or revert to internal knowledge. We chose this focus deliberately: classification enables objective evaluation on widely used datasets (e.g., e-snli), establishes strong baselines for comparability, and reflects many real-world applications such as content moderation, diagnosis, or categorization. At the same time, our analysis of generated explanations already shows similar domain- and definition-sensitivity patterns, suggesting that the core findings may extend beyond classification to generation and reasoning tasks, which we leave for future work.
To answer these questions systematically, we conduct experiments across four diverse datasets: e-snli for general natural language inference (Camburu et al., 2018), wellxplain for mental health (Garg, 2024), hatexplain for hate speech detection (Mathew et al., 2021), and wice for fact-checking (Kamoi et al., 2023). This multi-domain approach allows us to test whether our findings generalize beyond any single task or domain. Our main contributions are as follows: 1) We develop evaluation strategies to assess the performance and explainability of LLMs when provided with label definitions, which serve as a lightweight and interpretable form of external knowledge injection. 2) We design two probing strategies, definition permutation and definitional perturbation, to assess the sensitivity of LLMs to definition quality and alignment. 3) We conduct extensive experiments across four diverse benchmarks, e-snli, wellxplain, hatexplain, and wice, to assess model receptivity to label definitions.
@@ 2.1 Knowledge Conflict
Note. We distinguish these categories from disputed definitions, such as differing expert interpretations of terms like “fluency” or “readability.” These are not incorrect but represent legitimate divergences in perspective. While crucial in human evaluation settings, they introduce ambiguity that makes it difficult to isolate whether model performance changes arise from definition sensitivity or from inherent conceptual disagreement. For clarity, our study therefore focuses on clear right–wrong contrasts.
@@ 3 Setup and Methodology
To investigate definition receptivity across diverse model architectures, we selected four state-of-the-art LLMs with deliberately varied characteristics: GPT-4, LLaMA-3, Phi-3, and Mistral-7B.
This selection spans from compact 3.8B parameter models to trillion-parameter architectures, enabling us to determine whether investments in larger models yield proportional improvements in definition adherence—a key consideration for AI deployment in high-stakes domains where consistent interpretation of instructions is essential. Since definitions are concise, their inclusion adds minimal overhead in terms of token length. Nonetheless, this setup is primarily designed for short to medium-length inputs. For tasks requiring longer contexts (e.g., document-level inputs), the approach may not be suitable for smaller models like Phi-3, which has a 4k token limit in our case, thereby restricting the ability to integrate external definitions without truncation.
Each model represents distinct design philosophies with practical implications for deploying LLMs in definition-sensitive contexts: GPT-4 serves as our high-performance benchmark, representing proprietary models that organizations might access through APIs (Achiam et al., 2023); LLaMA-3 exemplifies open-source alternatives increasingly adopted in industry settings requiring customization and privacy (Meta, 2024); Phi-3 addresses the growing need for efficient, deployable models in resource-constrained environments (Bilenko, 2024); and Mistral-7B represents models optimized for the performance-efficiency trade-off crucial for commercial applications (Jiang et al., 2023). This diversity allows us to provide actionable insights on which architectural approaches best ensure definitional consistency across different deployment scenarios—whether in cloud infrastructure, edge devices, or hybrid settings. Detailed specifications for each model are provided in Table 5 (appendix˜A), with a representative prompt template shown in Table 6 (appendix˜A). For GPT-4, we balanced representation and cost by sampling 500 instances per label across datasets [^3] This approach ensures balanced class representation while managing computational resources, a practical constraint in production settings as well.. For all other models, we used complete test sets.
@@ 4 Experiments, Results, and Analysis
Model Permutation e-snli wellxplain hatexplain wice GPT-4 $\text{MCC}_{a}$ 0.83 0.35 0.42 0.32 $\overline{\text{MCC}}_{m}$ 0.74 0.17 0.27 0.16 LLaMA-3 $\text{MCC}_{a}$ 0.42 0.20 0.21 0.13 $\overline{\text{MCC}}_{m}$ 0.34 0.15 0.17 0.09 Phi-3 $\text{MCC}_{a}$ 0.24 0.32 0.20 0.05 $\overline{\text{MCC}}_{m}$ 0.23 0.11 0.06 0.04 Mistral-7B $\text{MCC}_{a}$ 0.51 0.10 0.05 0.05 $\overline{\text{MCC}}_{m}$ 0.48 0.07 0.05 0.03
@@
e-snli wellxplain hatexplain wice Model $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ $MCC_{Inc.}$ $MCC_{SInc.}$ $MCC_{Cor.}$ GPT-4 Meta-Response 0.83 0.83 0.31 0.31 0.38 0.31 0.36 0.44 0.13 0.31 0.32 LLaMA-3 0.08 0.32 0.36 0.16 0.20 0.21 0.10 0.13 0.15 0.09 0.13 0.13 Phi-3 0.22 0.21 0.23 0.29 0.35 0.38 0.11 0.08 0.14 0.04 0.01 0.05 Mistral-7B 0.20 0.48 0.40 0.18 0.20 0.20 0.05 0.07 0.09 0.03 0.03 0.05
Table 2: Impact of Definition Quality: Incorrect, Slightly Incorrect, and Correct Definitions across the general domain dataset (e-snli) and domain-specific datasets (wellxplain and hatexplain). Results highlight distinct patterns in model performance, with LLaMA-3 showing steady improvement from Incorrect to Correct, while Mistral-7B exhibits strong sensitivity to definition accuracy, particularly in domain-specific settings. Subscripts Inc., SInc., and Cor. denote Incorrect, Slightly Incorrect, and Correct, respectively. “Meta-Response” reflects the model declining to answer due to conflicting definitions.
@@ 4.1 LLM Receptivity: Results and Analysis
In the definition permutation scenario, Table 1 shows clear performance differences between aligned and misaligned label-definition mappings. Across all datasets, models achieve significantly higher MCC scores with correctly permuted definitions (aligned), with Mistral-7B on hatexplain being the striking exception where alignment produced minimal improvement—a surprising deviation from the otherwise consistent pattern. These findings align with the distributional hypothesis of language learning (Firth, 1957; Saxena et al., 2024), whereby LLMs learn word meanings through their contextual associations. When faced with conflicting cues—explicit definitions versus internalized patterns—models generally prioritize the explicit definitions provided in the prompt context, suggesting that instruction-following capabilities can override pre-trained associations.
The label definition accuracy scenario, summarized in Table 2, reveals varying sensitivities to definition quality across models and tasks. For e-snli, LLaMA-3 demonstrates remarkably strong responsiveness, with MCC values increasing more than threefold when moving from incorrect to slightly incorrect definitions, and nearly fourfold when comparing incorrect to correct definitions. Mistral-7B shows substantial but less dramatic improvements (140% and 100% respectively), while Phi-3 exhibits a surprisingly minimal sensitivity with near-negligible changes (-4.55% and +4.55%)—an unexpected resilience to definitional variations. Most unexpectedly, GPT-4 surprised us entirely by detecting definitional inconsistencies and declining to provide predictions—a sophisticated meta-cognitive response none of the other models exhibited. This pattern reveals a key distinction between linguistic competence (internalized knowledge of language structures) and performance (how that knowledge is applied in specific contexts), with larger models demonstrating stronger metacognitive capabilities to detect contradictions between provided definitions and their pre-training.
@@
For the wellxplain mental health categorization task, all models show moderate improvements with better definitions, though the sensitivity varies. GPT-4 and LLaMA-3 show similar improvement patterns (22.58% and 25% respectively) when moving from incorrect to more accurate definitions. Phi-3 demonstrates greater definitional sensitivity, with performance improving by approximately one-fifth when using slightly incorrect definitions and by nearly one-third with correct definitions. Contrary to expectations, Mistral-7B shows the least variation (11.11%), suggesting it relies more heavily on internal knowledge for this domain—a surprising finding given its stronger responsiveness in the e-SNLI task. This aligns with theories of knowledge acquisition, suggesting that when LLMs lack extensive pre-training exposure to specialized domains, they become more dependent on explicit contextual definitions, similar to how humans learn novel concepts by relying on explicit instruction when they lack prior experience.
In the hatexplain content moderation task, definition quality impacts performance substantially across all models. GPT-4’s accuracy improves steadily as definitions become more precise (16.13% with slightly incorrect and 41.94% with correct definitions). LLaMA-3 follows a similar pattern with more pronounced gains (30% and 50%). Interestingly, both Phi-3 and Mistral-7B show unexpectedly large improvements when given correct definitions (75% and 80% respectively)—a stark contrast to their behavior on other tasks. This suggests a “knowledge integration threshold” where models shift from relying on internal representations to trusting external definitions when the latter are sufficiently precise and consistent with their training.
For the wice fact-checking task, model performance under varying definition accuracies reveals minimal responsiveness across all models. Unlike other datasets, the MCC scores remain nearly flat, with negligible gains when moving from incorrect to correct definitions. LLaMA-3 and Mistral-7B show marginal improvement (from 0.09 to 0.13 and from 0.03 to 0.05, respectively), while Phi-3 even declines slightly from slightly incorrect to correct. These results suggest that wice, as an evidence-retrieval task, is less sensitive to semantic label definitions and more reliant on factual pattern matching—highlighting a limitation in current LLMs’ ability to incorporate abstract definitional guidance into decision-making for fact-based inference tasks.
@@
e-snli wellxplain hatexplain wice Model Scenario MCCMCC RougeRouge MCCMCC RougeRouge MCCMCC RougeRouge MCCMCC $F1$ Vanilla 0.86 0.21 0.41 0.06 0.33 0.05 0.28 0.71 GPT-4 Fixed Definition 0.80 0.20 0.46 0.05 0.36 0.04 0.32 0.73 Adjusted Definition 0.83 0.20 0.55 0.07 0.33 0.06 0.35 0.69 Fixed Definition+Few-shot 0.70 0.26 0.46 0.10 0.30 0.31 0.32 0.73 Vanilla 0.47 0.18 0.29 0.04 0.14 0.31 0.14 0.69 LLaMA-3 Fixed Definition 0.35 0.15 0.16 0.04 0.23 0.18 0.13 0.69 Adjusted Definition 0.36 0.14 0.44 0.40 0.14 0.10 0.14 0.69 Fixed Definition+Few-shot 0.36 0.20 0.38 0.38 0.15 0.04 0.11 0.74 Vanilla 0.48 0.19 0.33 0.06 0.13 0.01 0.03 0.68 Phi-3 Fixed Definition 0.38 0.12 0.38 0.06 0.19 0.02 0.05 0.68 Adjusted Definition 0.46 0.15 0.49 0.14 0.17 0.01 0.12 0.67 Fixed Definition+Few-shot 0.31 0.22 0.62 0.13 0.16 0.10 0.02 0.69 Vanilla 0.57 0.17 0.20 0.07 0.01 0.47 0.03 0.70 Mistral-7B Fixed Definition 0.32 0.20 0.20 0.07 0.12 0.35 0.05 0.69 Adjusted Definition 0.34 0.19 0.42 0.39 0.14 0.01 0.04 0.69 Fixed Definition+Few-shot 0.49 0.24 0.39 0.36 0.14 0.05 0.01 0.70
Table 3: When external guidance is actually useful: definitions make general NLI harder, but they are very helpful in specialized domains, and adjusted definitions also improve performance in wice for most models.
@@ 4.2 Definitions Integration: Results and Analysis
For e-snli, models perform best with the definition-free vanilla setting—a counterintuitive finding that challenges conventional wisdom about explicit guidance. While GPT-4 experiences modest degradation with definitions (6.98% and 3.49% decreases), LLaMA-3 and Mistral-7B show more substantial performance drops (23.40% to 43.86%). This pattern aligns with the bayesian inference framework of in-context learning proposed by Xie et al. (2021), where models already possess robust internal representations of common linguistic tasks through pretraining and may experience interference when explicit definitions contradict these learned representations (Min et al., 2022). Most surprisingly, adding few-shot examples alongside definitions further reduces performance across all models, with GPT-4 dropping by 18.60%—suggesting that exemplars can actually create confusion when combined with abstract definitions for general language tasks.